# Supplementary material – Approximate Dynamic Programming for Constrained Piecewise Affine Systems with Stability and Safety Guarantees

### v2

@@ I-A Backgrounds
There has been an increasing interest in control of piecewise affine (PWA) systems due to their capability of representing hybrid models and approximating nonlinear dynamics [1]. Many practical control problems can be modelled as PWA systems with constraints, including emergency evasive maneuvers [2], robotic manipulation that has multi-contact behaviors [3], and traffic control [4]. PWA systems are a special class of switched systems where the subsystem in each mode is affine and the transitions are based on the state belonging to different regions. Tractable controller design methods for PWA systems include synthesizing piecewise linear control laws via Linear Matrix Inequalities [5], adaptive control [6], and model predictive control (MPC) [5]. For suboptimal control of PWA systems with constraints, MPC is widely applied. However, MPC for PWA systems still faces challenges in computational complexity [7], because it involves solving a mixed-integer linear programming (MILP) problem. The complexity of solving MILP MPC problems is in general dominated by the number of integers, which is proportional to the prediction horizon. Explicit MPC [7], an offline version of MPC, requires solving a parametric MILP problem, which is also suffering from computational complexity issues. These issues make MPC only suitable for slow PWA processes or for low-dimensional problems with short prediction horizons [3].
In contrast to MPC, reinforcement learning (RL) can learn a policy that minimizes a finite-/infinite-horizon cost and has a much lower online computational burden [8] than MPC. In RL, two different methodologies can be distinguished: policy search and dynamic programming [9]. Policy search directly searches for an optimal policy, while dynamic programming aims to solve the Bellman equation first and then determines the policy. Dynamic programming has the advantage over policy search is that it reduces the policy optimization problem to a one-step look-ahead problem. When applied to systems with continuous state and input spaces, approximate dynamic programming (ADP), which uses sample-based methods accompanied by sophisticated function approximators, has been developed [10, 11]. In this paper, we consider approximate value iteration (VI), the most basic and direct way to solve the Bellman equation. Moreover, we provide comprehensive performance guarantees for stability, safety and sub-optimality of the developed ADP approach.
@@ I-B Related work
Approximate MPC: Approximate MPC parameterizes a policy and then uses supervised learning or gradient-based methods to mimic a predictive control policy. The online computational cost is thus significantly reduced because approximate MPC directly computes control actions based on the learned parameters, rather than online solving an optimization problem. Some approximate MPC work considers linear systems, with different focuses on, e.g., stability [12] and constraint satisfaction [13]. Some approximate MPC approaches can handle nonlinear control problems with constraints, e.g., by using constraint tightening [14]. However, as these approaches are developed for general nonlinear systems, they lack deterministic performance verification tools. RL for constrained control: RL for constrained control can be roughly categorized into two groups: policy-projection RL methods, which impose constraints after learning, and policy-optimization RL methods, which impose constraints during learning. Using RL for controller synthesis allows designers to stop the learning process and get a satisfactory policy whenever the expected performance is met. For the first group, a predictive safety filter [15, 16] can be adopted to modify the learned policy, which can be derived from any RL algorithm. This group of approaches can in general be applied to constrained nonlinear control problems. For PWA systems with linear constraints, this group needs to solve mixed-integer convex problems online [16], and as a result it is not suitable for large-scale systems or for systems requiring fast computation. Additionally, this group of methods in general jeopardizes optimality and stability [17]. For the second group, constrained policy optimization [18, 19, 20] is often used. Most methods [21] consider constrained Markov Decision Process (MDP) problems, in which constraints are on expected cumulative costs. Recent developments have been made to transform instantaneous constraints into constraints on expected cumulative costs. Among these developments, [22] considers instantaneous expectation constraints, while [23] contemplates chance constraints. In [24], hard instantaneous constraints are considered, but the convergence of the RL algorithm relies on the assumptions of the finiteness of MDP and the continuous differentiability of the value function, which is not the case for optimal control of PWA systems. Besides, [24, 22] do not provide stability or safety analysis of the RL policy. Lyapunov stable RL: In addition to the online computational advantage, the stability property of RL controllers has been investigated recently. In [25], stability analysis of approximate VI for unconstrained nonlinear systems is conducted, but it results in very restrictive conditions, which require the approximation errors per iteration to be upper bounded. [26, 27] further relax these conditions by only focusing on the stopping criterion and the approximation error in the last iteration. Nevertheless, these references consider unconstrained problems and do not address the sub-optimality of the RL policy. Besides, the combination of RL and MPC is proposed in [28, 27, 29, 30] as a theoretically justified methodology to produce safe and stabilizing policies. However, in [28, 27, 29, 30], the online computational issue of MPC is not solved by introducing RL. Based on these observations, for the policy optimization methods, one could conclude that (i) no work has been done yet for PWA systems, and (ii) good and comprehensive performance of RL-based controllers regarding online computing convenience, stability, and constraint fulfillment cannot be achieved simultaneously. In this paper, we will consider policy optimization methods for PWA systems subject to linear or union-of-polyhedra (UoP) constraints, and solve the above two issues. Meanwhile, most RL/ADP methods ignore the properties of the value function and the policy, e.g., boundedness, continuity, nonlinearity, so that the design of function approximators is usually done in an empirical way [31]. For PWA systems with a linear performance index, fortunately, extant research on explicit hybrid MPC has shown that the optimal value function and the optimal policy are both PWA [32, 1]. This provides an opportunity to use PWA function approximation, as will be investigated in this paper.
@@ III Value Iteration with Penalty Functions
${J^{\mathrm{soft}*}}\left({{x}}\right)=\mathop{\min}\limits_{\{u_{i},x_{i}\}^{\infty}_{i=0}}$ $\sum_{i=0}^{\infty}l_{p}\left(x_{i},u_{i}\right)+P(x_{i},X)$
[…]
@@ V-A Stability and safety analysis
Consider Algorithm 1. Suppose that Assumptions 1-2 hold in option (a) and Assumptions 1-3 hold in option (b). (i) Then, there exists a positive constant $\gamma<\infty$ such that for all $k\geq 0$, $J_{k}^{\mathrm{soft}}(x)\leq\gamma l(x,0)$, $\forall x\in\bar{X}$; (ii) there exists a finite $\bar{k}>0$ such that for all $k\geq\bar{k}$, we have
@@
Consider Algorithm 2 and the proposed policies $\hat{\pi}^{\mathrm{im}}(\cdot)$ and $\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}(\cdot)$. Let $\Omega$ be a compact subset of $X$. Suppose that Assumptions 1-2 hold in option (a) and Assumptions 1-3 hold in option (b). Consider the follow conditions: (C1): There exist a constant $\zeta\in(0,1)$ and a positive integer $k$ such that $\left|{{{{{\hat{J}}_{k-1}}(x)-J_{k-1}^{\mathrm{soft}}(x)}}}\right|\leq\zeta{{J_{k-1}^{\mathrm{soft}}(x)}},\;\forall x\in\Omega$. (C2): There exist a constant $e_{\mathrm{p}}>0$ and a positive integer $k$ such that $\hat{J}_{\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}}(x)-\hat{J}_{\hat{\pi}^{\mathrm{im}}}(x)\leq e_{\mathrm{p}}l(x,0),\;\forall x\in\Omega$. Here, $\hat{J}_{\pi}(\cdot)$ is defined by $\hat{J}_{\pi}(x)\triangleq l(x,\pi(x))+\hat{J}_{k-1}(f_{\mathrm{PWA}}(x,\pi(x)))$. As a result, we have: (i) If C1 holds with $k\geq\bar{k}$ and
@@
where $\bar{k}$, $\beta$, and $\gamma$ come from Lemma 2, the closed-loop system $x_{t+1}=f_{\mathrm{PWA}}(x_{t},\hat{\pi}^{\mathrm{im}}(x_{t})),\;t=0,1,...$ is asymptotically stable and safe in $\mathcal{B}(J^{\mathrm{soft}}_{k-1},\Omega)\cap\mathcal{B}(\hat{J}_{k-1},\Omega)$. (ii) If C1 and C2 hold with $k\geq\bar{k}$, and
@@
Although Theorem 2 provides sufficient conditions for stability, some of them (e.g., C1 and C2) are difficult to verify, unless the explicit form of $J_{k-1}^{\mathrm{soft}}(\cdot)$ is known. A method that can verify C1 and C2 in a probabilistic way is reported in [14]. On the other hand, Theorem 2 suggests several ways to design the parameters and function approximators in Algorithm 2.
@@ V-B Sub-optimality analysis
Consider Algorithm 2 and the proposed policies $\hat{\pi}^{\mathrm{im}}(\cdot)$ and $\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}(\cdot)$. Let $\Omega$ be a compact subset of $X$. (i) Let the assumptions in (i) of Theorem 2 hold and $12\zeta\gamma<$ 1. Then, for any $x\in\mathcal{B}(J^{\mathrm{soft}}_{k-1},\Omega)\cap\mathcal{B}(\hat{J}_{k-1},\Omega)$, we have the inequality
@@
(22) and (23) confirm the intuition that a smaller approximation error of the critic (and the actor) leads to tighter sub-optimality guarantees. Namely, as $\zeta\to 0,\;e_{\mathrm{p}}\to 0$, and $k\to\infty$, we have $J_{\hat{\pi}^{\mathrm{im}}}(x)\to J^{\mathrm{soft*}}(x)$ and $J_{\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}}(x)\to J^{\mathrm{soft*}}(x)$ for any $x\in\mathcal{B}(J^{\mathrm{soft}}_{k-1},\Omega)\cap\mathcal{B}(\hat{J}_{k-1},\Omega)$. Moreover, if option (a) of Algorithm 2 is used, $J^{\mathrm{soft}}_{k-1}(x)$ in (22) and (23) can be replaced by $J^{\mathrm{soft*}}(x)$ because $J^{\mathrm{soft}}_{k-1}(x)\leq J^{\mathrm{soft*}}(x),\;\forall x\in\mathcal{X}$.
@@ VI-A Example 1: inverted pendulum with elastic walls
(b) Closed-loop trajectories of the inverted pendulum. The shaded region is $\bar{X}$. The curves with different colors represent the state trajectories starting from different vertices.
(c) The learning process of the ADP algorithm with different options.
(d) Safe and stable regions verified by the verification framework.
(e) Statistical analysis of different methods regarding CPU times. D-max means the difference of two max-affine functions.
(f) Closed-loop trajectories of the inverted pendulum with UoP state constraints. The shaded region is $\bar{X}$. The curves with different colors represent the state trajectories starting from different vertices.
@@
To carry out option (b) of Algorithm 2, we consider the linear subsystem $(A_{3},\;B_{3})$ and readily get a linear stabilizing feedback law $u_{0}=[-40\;\;-10]x$, which is used to initialize the value function of option (b). The optimal value function $J^{*}(\cdot)$ is computed by the MPT3 toolbox [49] in 9.6 hours. The learning processes of different versions of Algorithm 2 are compared in Fig. 2(c), in which the mean square errors between $\hat{J}_{k-1}\left(x^{(i)},\theta_{k}\right)$ and $J^{*}(x^{(i)})$ are depicted. One can see that the value function approximation in option (b) has a much faster convergence rate than that in option (a), and option 2 results in lower approximation errors than option 1. The accelerated convergence rate in option (b) is mainly attributed to its initial value function approximation $\hat{J}_{0}\left(\cdot,\theta_{0}\right)$ being considerably closer to $J^{\mathrm{soft}*}(\cdot)$ compared to option (a). The ultimately reduced approximation errors in option 2 likely stem from a more densely sampled state space.
@@
The conditions for the stability and safety of the policy $\hat{\pi}_{\mathrm{proj}}^{\mathrm{ex}}(\cdot)$ are verified in Fig. 2(d). Fig. 2(d) depicts the safe and stable regions that are analyzed by the proposed verification framework. The input constraints are always satisfied in the simulation because the optimization problems (14) and (17) contain the input constraints as hard constraints. The blue region represents $\mathcal{B}_{r_{1}}$, which is computed by (V-C) with $r_{1}=18,\;c_{1}=0.1,\;r_{2}=3$. We note that for some states in $\mathcal{B}_{r_{2}}$, which is colored red, the objective function in (V-C) becomes positive, so the verification method in [34] fails. However, (V-C) outputs a negative $a^{*}_{2}$ with $c_{2}=0.1$, which means that any trajectory of the closed-loop system starting from $\mathcal{B}_{r_{1}}$ will reach in the neighborhood $\mathcal{B}_{r_{2}}$ containing the origin in finite number of time steps. Furthermore, the safe and stable region $\mathcal{B}_{r_{1}}$ is enlarged to $X_{\mathrm{in}}$ (the yellow region) by (V-C) with $N=3$. However, we observe that the verified safe and stable polytope $X_{\mathrm{in}}$ may be conservative compared to our trajectory simulation in Fig. 2(b). Suppose that $X_{\mathrm{in}}\triangleq\{x\in\mathbb{R}^{n_{x}}|E_{X_{\mathrm{in}}}x\leq g_{X_{\mathrm{in}}}\}$. The conservatism is primarily attributed to the naive choice of $E_{X_{\mathrm{in}}}$. Practical strategies to mitigate this conservatism involve using a UoP $X_{\mathrm{in}}$ and extending the horizon $N$. These two methods, on the other hand, will increase the complexity of the MILP problem (V-C).
@@
We compare the CPU time of running the ADP-based controllers, hybrid MPC, explicit MPC computed by the MPT3 toolbox, nonlinear MPC that uses the nonlinear model, action-governor RL in [16], and the predictive safety filter (PSF) in [15]. With an MPC horizon 8, the implementation of hybrid MPC requires to solve an MILP problem that contains 94 continuous variables, 28 binary variables, 328 inequality constraints, and 21 equality constraints. In comparison, for $\hat{\pi}^{\mathrm{im}}(\cdot)$ with the value function approximator chosen as the ReLU NN (or the difference of two max-affine functions that has 10 and 3 terms in the first and second max blocks), the resulting MILP problem has 29(15) continuous variables, 20(7) binary variables, 94(73) inequality constraints, and 3(4) equality constraints. We only compare the time of computation that needs the knowledge of the current state, including constructing and solving optimization problems. For statistical analysis, we randomly select 100 initial states and run the system in 50 time steps [^2] The number of time steps is set based on the fact that in our numerical experiments most scenarios converge to the origin after this number of time steps.. The average computation time per time step of the different methods is shown in Fig. 2(e). The computation of $\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}(\cdot)$ requires the least amount of CPU time, which is around $2.2\times 10^{-6}$ s per time step. Besides, $\hat{\pi}^{\mathrm{im}}(\cdot)$ performs better than hybrid MPC, explicit MPC, action-governor RL, and PSF, regarding online computation time, both when using the ReLU NN (about 0.028 s) or the difference of two max-affine functions (about 0.026 s) for value function approximation.

### v3

@@ I-A Backgrounds
There has been an increasing interest in control of piecewise affine (PWA) systems due to their capability of representing hybrid models and approximating nonlinear dynamics [1]. Many practical control problems can be modelled as PWA systems with constraints, including emergency evasive maneuvers [2], robotic manipulation that has multi-contact behaviors [3], and traffic control [4]. PWA systems are a special class of switched systems where the subsystem in each mode is affine and the transitions are based on the state belonging to different regions. Tractable controller design methods for PWA systems include synthesizing piecewise linear control laws via Linear Matrix Inequalities [5], adaptive control [6], and model predictive control (MPC) [5]. The challenges of controlling a PWA system include ensuring stability and achieving optimality guarantees [5], as well as addressing both offline and online computational complexity [7, 3]. These difficulties primarily arise from the system’s hybrid structure and inherent nonlinearity. For suboptimal control of PWA systems with constraints, MPC is widely applied. However, MPC for PWA systems still faces challenges in computational complexity [7], because it involves solving a mixed-integer linear programming (MILP) problem. The complexity of solving MILP MPC problems is in general dominated by the number of integers, which is proportional to the prediction horizon. Explicit MPC [7], an offline version of MPC, requires solving a parametric MILP problem, which is also suffering from computational complexity issues. These issues make MPC only suitable for slow PWA processes or for low-dimensional problems with short prediction horizons [3].
In contrast to MPC, reinforcement learning (RL) can learn a policy that minimizes a finite-/infinite-horizon cost and has a much lower online computational burden [8] than MPC. In RL, two different methodologies can be distinguished: policy search and dynamic programming [9]. Policy search directly searches for an optimal policy, while dynamic programming aims to solve the Bellman equation first and then determines the policy. Dynamic programming has the advantage over policy search is that it reduces the policy optimization problem to a one-step look-ahead problem. When applied to systems with continuous state and input spaces, approximate dynamic programming (ADP), which uses sample-based methods accompanied by sophisticated function approximators, has been developed [10, 11]. In this paper, we consider approximate value iteration (VI), the most basic and direct way to solve the Bellman equation. Moreover, we provide comprehensive performance guarantees for stability, safety and sub-optimality of the developed ADP approach. In the context of RL, safety can have various definitions. In this paper, we specifically address safety as ensuring that the state and input of the system satisfy predefined constraints throughout the entire system’s evolution after the learning process is finished.
@@ I-B Related work
Approximate MPC: Approximate MPC parameterizes a policy and then uses supervised learning or gradient-based methods to mimic a predictive control policy. The online computational cost is thus significantly reduced because approximate MPC directly computes control actions based on the learned parameters, rather than online solving an optimization problem. Some approximate MPC work considers linear systems, with different focuses on, e.g., stability [12] and constraint satisfaction [13]. Some approximate MPC approaches can handle nonlinear control problems with constraints, e.g., by using constraint tightening [14]. However, as these approaches are developed for general nonlinear systems, they lack deterministic performance verification tools.
RL for constrained control: RL for constrained control can be roughly categorized into two groups: policy-projection RL methods, which impose constraints after learning, and policy-optimization RL methods, which impose constraints during learning. Using RL for controller synthesis allows designers to stop the learning process and get a satisfactory policy whenever the expected performance is met.
For the first group, a predictive safety filter [15, 16] can be adopted to modify the learned policy, which can be derived from any RL algorithm. This group of approaches can in general be applied to constrained nonlinear control problems. For PWA systems with linear constraints, this group needs to solve mixed-integer convex problems online [16], and as a result it is not suitable for large-scale systems or for systems requiring fast computation. Additionally, this group of methods in general jeopardizes optimality and stability [17].
For the second group, constrained policy optimization [18, 19, 20] is often used. Most methods [21] consider constrained Markov Decision Process (MDP) problems, in which constraints are on expected cumulative costs. Recent developments have been made to transform instantaneous constraints into constraints on expected cumulative costs. Among these developments, [22] considers instantaneous expectation constraints, while [23] contemplates chance constraints. In [24], hard instantaneous constraints are considered, but the convergence of the RL algorithm relies on the assumptions of the finiteness of MDP and the continuous differentiability of the value function, which is not the case for optimal control of PWA systems. Besides, [24, 22] do not provide stability or safety analysis of the RL policy.
Lyapunov stable RL: In addition to the online computational advantage, the stability property of RL controllers has been investigated recently. In [25], stability analysis of approximate VI for unconstrained nonlinear systems is conducted, but it results in very restrictive conditions, which require the approximation errors per iteration to be upper bounded. [26, 27] further relax these conditions by only focusing on the stopping criterion and the approximation error in the last iteration. Nevertheless, these references consider unconstrained problems and do not address the sub-optimality of the RL policy. Besides, the combination of RL and MPC is proposed in [28, 27, 29, 30] as a theoretically justified methodology to produce safe and stabilizing policies. However, in [28, 27, 29, 30], the online computational issue of MPC is not solved by introducing RL.
Based on these observations, for the policy optimization methods, one could conclude that (i) no work has been done yet for PWA systems, and (ii) good and comprehensive performance of RL-based controllers regarding online computing convenience, stability, and constraint fulfillment cannot be achieved simultaneously. In this paper, we will consider policy optimization methods for PWA systems subject to linear or union-of-polyhedra (UoP) constraints, and solve the above two issues.
Meanwhile, most RL/ADP methods ignore the properties of the value function and the policy, e.g., boundedness, continuity, nonlinearity, so that the design of function approximators is usually done in an empirical way [31]. For PWA systems with a linear performance index, fortunately, extant research on explicit hybrid MPC has shown that the optimal value function and the optimal policy are both PWA [32, 1]. This provides an opportunity to use PWA function approximation, as will be investigated in this paper.
@@ III Value Iteration with Penalty Functions
${J^{\mathrm{soft}*}}\left({{x}}\right)=\mathop{\min}\limits_{\{u_{i},x_{i}\}^{\infty}_{i=0}}$ $\sum_{i=0}^{\infty}l_{\mathrm{p}}\left(x_{i},u_{i}\right)+P(x_{i},X)$
$\mathrm{s.t.}\;\;\;\;$ ${x_{i+1}}={f_{{\text{PWA}}}}({x_{i}},{u_{i}}),\;u_{i}\in U,\;i=0,1,...$
@@
After the optimal value function $J_{k}^{\text{soft }}$ is obtained by Algorithm 1, the control policy is implicitly determined by the solution of the optimization problem (II-A) with $J^{*}$ replaced by $J_{k}^{\text{soft }}$.
Remark 1
@@ V-A Stability and safety analysis
Consider Algorithm 1. Suppose that Assumptions 1-2 hold in option (a) and Assumptions 1-3 hold in option (b).
(i) Then, there exists a positive constant $\gamma<\infty$ such that for all $k\geq 0$, $J_{k}^{\mathrm{soft}}(x)\leq\gamma l(x,0)$, $\forall x\in\bar{X}$;
(ii) there exists a finite $\bar{k}>0$ such that for all $k\geq\bar{k}$, we have
@@
Consider Algorithm 2 and the proposed policies $\hat{\pi}^{\mathrm{im}}(\cdot)$ and $\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}(\cdot)$. Let $\Omega$ be a compact subset of $X$. Suppose that Assumptions 1-2 hold in option (a) and Assumptions 1-3 hold in option (b). Consider the follow conditions:
(C1): There exist a constant $\zeta\in(0,1)$ and a positive integer $k$ such that $\left|{{{{{\hat{J}}_{k-1}}(x)-J_{k-1}^{\mathrm{soft}}(x)}}}\right|\leq\zeta{{J_{k-1}^{\mathrm{soft}}(x)}},\;\forall x\in\Omega$.
(C2): There exist a constant $e_{\mathrm{p}}>0$ and a positive integer $k$ such that $\hat{J}_{\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}}(x)-\hat{J}_{\hat{\pi}^{\mathrm{im}}}(x)\leq e_{\mathrm{p}}l(x,0),\;\forall x\in\Omega$. Here, $\hat{J}_{\pi}(\cdot)$ is defined by $\hat{J}_{\pi}(x)\triangleq l(x,\pi(x))+\hat{J}_{k-1}(f_{\mathrm{PWA}}(x,\pi(x)))$.
As a result, we have:
(i) If C1 holds with $k\geq\bar{k}$ and
@@
where $\bar{k}$, $\beta$, and $\gamma$ come from Lemma 2, the closed-loop system $x_{t+1}=f_{\mathrm{PWA}}(x_{t},\hat{\pi}^{\mathrm{im}}(x_{t})),\;t=0,1,...$ is asymptotically stable and safe in $\mathcal{B}(J^{\mathrm{soft}}_{k-1},\Omega)\cap\mathcal{B}(\hat{J}_{k-1},\Omega)$.
(ii) If C1 and C2 hold with $k\geq\bar{k}$, and
@@
Remark 4
Although Theorem 2 provides sufficient conditions for stability, some of them (e.g., C1 and C2) are difficult to verify. A method that can verify C1 and C2 in a probabilistic way is reported in [14]. On the other hand, Theorem 2 suggests several ways to design the parameters and function approximators in Algorithm 2. For practical verification of stability and safety, in Section V.C, we move beyond using Conditions C1 and C2. Instead, we propose an offline verification framework based on solving MILP problems. As will be demonstrated in the case study, this framework effectively verifies safe and stable regions.
@@ V-B Sub-optimality analysis
Consider Algorithm 2 and the proposed policies $\hat{\pi}^{\mathrm{im}}(\cdot)$ and $\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}(\cdot)$. Let $\Omega$ be a compact subset of $X$.
(i) Let the assumptions in (i) of Theorem 2 hold and $12\zeta\gamma<$ 1. Then, for any $x\in\mathcal{B}(J^{\mathrm{soft}}_{k-1},\Omega)\cap\mathcal{B}(\hat{J}_{k-1},\Omega)$, we have the inequality
@@
It is almost impossible to get an optimal control policy from (IV-B) due to the approximation error. However, (22) and (23) confirm the intuition that a smaller approximation error of the critic (and the actor) leads to tighter sub-optimality guarantees. Namely, as $\zeta\to 0,\;e_{\mathrm{p}}\to 0$, and $k\to\infty$, we have $J_{\hat{\pi}^{\mathrm{im}}}(x)\to J^{\mathrm{soft*}}(x)$ and $J_{\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}}(x)\to J^{\mathrm{soft*}}(x)$ for any $x\in\mathcal{B}(J^{\mathrm{soft}}_{k-1},\Omega)\cap\mathcal{B}(\hat{J}_{k-1},\Omega)$. Moreover, if option (a) of Algorithm 2 is used, $J^{\mathrm{soft}}_{k-1}(x)$ in (22) and (23) can be replaced by $J^{\mathrm{soft*}}(x)$ because $J^{\mathrm{soft}}_{k-1}(x)\leq J^{\mathrm{soft*}}(x),\;\forall x\in\mathcal{X}$.
@@ VI-A Example 1: inverted pendulum with elastic walls
(b) The learning process of the ADP algorithm with different options.
(c) Closed-loop trajectories of the inverted pendulum. The shaded region is $\bar{X}$.
(d) Time-domain system responses.
(e) Safe and stable regions verified by the verification framework.
(f) Statistical analysis of different methods regarding CPU times. D-max means the difference of two max-affine functions.
(g) Closed-loop trajectories of the inverted pendulum with UoP state constraints.
(h) Time-domain system responses under UoP state constraints.
@@
To carry out option (b) of Algorithm 2, we consider the linear subsystem $(A_{3},\;B_{3})$ and readily get a linear stabilizing feedback law $u_{0}=[-40\;\;-10]x$, which is used to initialize the value function of option (b). The optimal value function $J^{*}(\cdot)$ is computed by the MPT3 toolbox [49] in 9.6 hours. The learning processes of different versions of Algorithm 2 are compared in Fig. 2(b), in which the mean square errors between $\hat{J}_{k-1}\left(x^{(i)},\theta_{k}\right)$ and $J^{*}(x^{(i)})$ are depicted. One can see that the value function approximation in option (b) has a much faster convergence rate than that in option (a), and option 2 results in lower approximation errors than option 1. The accelerated convergence rate in option (b) is mainly attributed to its initial value function approximation $\hat{J}_{0}\left(\cdot,\theta_{0}\right)$ being considerably closer to $J^{\mathrm{soft}*}(\cdot)$ compared to option (a). The ultimately reduced approximation errors in option 2 likely stem from a more densely sampled state space.
@@
The conditions for the stability and safety of the policy $\hat{\pi}_{\mathrm{proj}}^{\mathrm{ex}}(\cdot)$ are verified in Fig. 2(e). Fig. 2(e) depicts the safe and stable regions that are analyzed by the proposed verification framework. The input constraints are always satisfied in the simulation because the optimization problems (14) and (17) contain the input constraints as hard constraints. The blue region represents $\mathcal{B}_{r_{1}}$, which is computed by (V-C) with $r_{1}=18,\;c_{1}=0.1,\;r_{2}=3$. We note that for some states in $\mathcal{B}_{r_{2}}$, which is colored red, the objective function in (V-C) becomes positive, so the verification method in [34] fails. However, (V-C) outputs a negative $a^{*}_{2}$ with $c_{2}=0.1$, which means that any trajectory of the closed-loop system starting from $\mathcal{B}_{r_{1}}$ will reach in the neighborhood $\mathcal{B}_{r_{2}}$ containing the origin in finite number of time steps. Furthermore, the safe and stable region $\mathcal{B}_{r_{1}}$ is enlarged to $X_{\mathrm{in}}$ (the yellow region) by (V-C) with $N=3$. However, we observe that the verified safe and stable polytope $X_{\mathrm{in}}$ may be conservative compared to our trajectory simulation in Fig. 2(c). Suppose that $X_{\mathrm{in}}\triangleq\{x\in\mathbb{R}^{n_{x}}|E_{X_{\mathrm{in}}}x\leq g_{X_{\mathrm{in}}}\}$. The conservatism is primarily attributed to the naive choice of $E_{X_{\mathrm{in}}}$. Practical strategies to mitigate this conservatism involve using a UoP $X_{\mathrm{in}}$ and extending the horizon $N$. These two methods, on the other hand, will increase the complexity of the MILP problem (V-C).
@@
We compare the CPU time of running the ADP-based controllers, hybrid MPC, explicit MPC computed by the MPT3 toolbox, nonlinear MPC that uses the nonlinear model, action-governor RL in [16], and the predictive safety filter (PSF) in [15]. With an MPC horizon 8, the implementation of hybrid MPC requires to solve an MILP problem that contains 94 continuous variables, 28 binary variables, 328 inequality constraints, and 21 equality constraints. In comparison, for $\hat{\pi}^{\mathrm{im}}(\cdot)$ with the value function approximator chosen as the ReLU NN (or the difference of two max-affine functions that has 10 and 3 terms in the first and second max blocks), the resulting MILP problem has 29(15) continuous variables, 20(7) binary variables, 94(73) inequality constraints, and 3(4) equality constraints. We only compare the time of computation that needs the knowledge of the current state, including constructing and solving optimization problems. For statistical analysis, we randomly select 100 initial states and run the system in 50 time steps [^3] The number of time steps is set based on the fact that in our numerical experiments most scenarios converge to the origin after this number of time steps.. The average computation time per time step of the different methods is shown in Fig. 2(f). The computation of $\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}(\cdot)$ requires the least amount of CPU time, which is around $2.2\times 10^{-6}$ s per time step. Besides, $\hat{\pi}^{\mathrm{im}}(\cdot)$ performs better than hybrid MPC, explicit MPC, action-governor RL, and PSF, regarding online computation time, both when using the ReLU NN (about 0.028 s) or the difference of two max-affine functions (about 0.026 s) for value function approximation.